The monthly amount spent on stationery in a large company has a normal distribution with mean R2500 and standard deviation R150. How much should the company budget for stationery if the chance of exceeding the budget has to be only 2.5%?


1
Expert's answer
2022-05-23T16:14:39-0400
"P(X>x)=1-P(Z\le \dfrac{x-2500}{150})=0.025"

"\dfrac{x-2500}{150}=1.96"

"x=2500+150(1.96)"

"x=R2794"